### v1

# Conditional Panoramic Image Generation via Masked Autoregressive Modeling

Abstract
Recent progress in panoramic image generation has underscored two critical limitations in existing approaches. First, most methods are built upon diffusion models, which are inherently ill-suited for equirectangular projection (ERP) panoramas due to the violation of the identically and independently distributed (i.i.d.) Gaussian noise assumption caused by their spherical mapping. Second, these methods often treat text-conditioned generation (text-to-panorama) and image-conditioned generation (panorama outpainting) as separate tasks, relying on distinct architectures and task-specific data. In this work, we propose a unified framework, Panoramic AutoRegressive model (PAR), which leverages masked autoregressive modeling to address these challenges. PAR avoids the i.i.d. assumption constraint and integrates text and image conditioning into a cohesive architecture, enabling seamless generation across tasks. To address the inherent discontinuity in existing generative models, we introduce circular padding to enhance spatial coherence and propose a consistency alignment strategy to improve generation quality. Extensive experiments demonstrate competitive performance in text-to-image generation and panorama outpainting tasks while showcasing promising scalability and generalization capabilities. Project Page: https://wang-chaoyang.github.io/project/par.
Figure 1: Generated samples from Panoramic AutoRegressive (PAR) Model. PAR unifies several conditional panoramic image generation tasks, including text-to-panorama, panorama outpainting, and panoramic image editing.
1 Introduction
Panoramic image generation[1, 11, 12, 17, 18] has recently emerged as a pivotal research direction in computer vision, driven by its capability to capture immersive 360-degree horizontal and 180-degree vertical fields of view (FoV). This unique visual representation closely mimics human panoramic perception, enabling transformative applications across VR/AR [33, 76], autonomous driving [42], visual navigation [5, 24], and robotics [43, 58].
Despite some progress in recent years, existing panorama generation methods [17, 54, 55, 71, 67, 26, 35, 59, 75, 62, 61, 74] predominantly rely on diffusion-based base models and face two significant problems: task separation and inherent limitations of the modeling approach. First, different panoramic generation tasks are highly bound to specific foundation models. For instance, most text-to-panorama models [17, 54, 55, 71, 67] are based on Stable Diffusion (SD) [46, 41] while panorama outpainting models [26, 35, 59] are fine-tuned on SD-inpainting variants, resulting in task-specific pipelines that lack flexibility. Second, diffusion-based models struggle to handle the unique characteristics of equirectangular projection (ERP) panoramas, which map spherical data onto a 2D plane, violating the i.i.d. Gaussian noise assumption.
In addition to these fundamental issues, redundant modeling in existing methods introduces unnecessary computational costs. This redundancy manifests in several ways: 1) Some approaches [28, 35] iteratively inpaint and warp local areas, leading to error accumulation over successive steps. 2) Others [71, 53] adopt a dual-branch structure, with global and local branches, where the local branches are not directly used as final output targets. These challenges highlight the lack of a simple, unified, and efficient paradigm for conditional panoramic image generation.
In this work, we explore a powerful yet much less explored panoramic image generation paradigm to address the above limitations - the violation of i.i.d. assumption and the lack of unification. Moreover, we allow the model to take text or images as input and control the generated content at any position on the canvas. As shown in Fig. 1, PAR is competent for text-to-panorama, panorama outpainting, and editing tasks. Notably, our PAR requires no task-specific data engineering - both conditions are seamlessly integrated via a single likelihood objective. In comparison, Omni2 [65] relies on meticulously designed joint training datasets to align heterogeneous tasks, inevitably introducing complexity and domain bias.
For this goal, inspired by masked autoregressive modeling (MAR) [29], we propose PAR. On the one hand, it is based on AR modeling, which avoids the constraints of i.i.d. assumptions. On the other hand, it allows for generation in arbitrary order, overcoming the shortcoming of traditional raster-scan methods that are incapable of image-condition generation. This flexibility of PAR unifies conditional generation tasks - panorama outpainting reduces to a subset of text-to-panorama synthesis, where existing image tokens serve as partially observed sequences. Building upon the MAR-based generative framework, we introduce two critical enhancements to address inherent limitations in standard architectures. First, conventional VAE-based compression often exhibits spatial bias during encoding-decoding: center regions benefit from contextual information across all directions, while peripheral pixels suffer from unidirectional receptive fields, contradicting the cyclic nature of ERP panoramas. To mitigate this, we propose dual-space circular padding, applying cyclic padding operations in both latent and pixel spaces. This ensures seamless feature propagation across ERP boundaries, aligning geometric and semantic continuity. Second, as the foundation model is pre-trained on massive perspective images, we devise a translation consistency loss. This loss explicitly regularizes the model to accommodate the prior of ERP panoramas by minimizing discrepancies between the model’s outputs under shifted and original inputs. Ultimately, two innovations holistically reconcile the unique demands of 360-degree imagery with the flexibility of MAR modeling. Experiments demonstrate that PAR outperforms specialist models in the text-to-panorama task, with 37.37 FID on the Matterport3D dataset. Moreover, on the outpainting task, it has better generation quality and avoids the problem of repeated structure generation. Ablation studies demonstrate the model’s scalability in terms of computational cost and parameters, as well as its generalization to out-of-distribution (OOD) data and other tasks, such as zero-shot image editing.
Our contribution can be summarized as follows: 1) We propose PAR, which fundamentally avoids the conflict between ERP and i.i.d. assumption raised by diffusion models, and also seamlessly integrates text and image-conditioned generation within a single architecture. 2) To enforce better adaptation to the panorama characteristic, we propose specialized designs, including dual-space circular padding and a translation consistency loss. 3) We evaluate our model on popular benchmarks, demonstrating competitive performance in text and image-conditional panoramic image generation tasks and showcasing its scalability and generalization capabilities.
2 Related Work
Conditional Panoramic Image Generation. There are two primary task paradigms in this area: 1) text-to-panorama (T2P), which synthesizes 360-degree scenes from textual descriptions, and 2) panorama outpainting (PO), which extends partial visual inputs to full panoramas. Given the success of diffusion models (DMs) [30, 31, 32, 41, 46, 49, 50] on generation tasks, such as image generation [46, 72, 22, 41, 40, 16], image editing [3, 36, 47, 73], image super resolution [21, 48, 68], video generation [2, 19, 66], current approaches typically build upon DMs, yet face two problems. First, DM is inherently unsuitable for panorama tasks due to reliance on the i.i.d. Gaussian noise assumption, which ERP transformations violate. Second, existing approaches lack task unification. Specifically, T2P methods typically fine-tune SD, while PO solutions adapt the SD-inpainting variants. Although some work [65] attempts a unified solution, it is based on DM and relies on complex task-specific data engineering. From a methodological perspective, existing methods can be divided into bottom-up and top-down. The former generates images with limited FoV sequentially [28, 35] or simultaneously [55, 26] and finally stitches them together, but faces the problem of lacking global information and error accumulation. The latter [71, 53] usually adopts a dual-branch architecture including a panorama and a perspective branch. However, this design brings unnecessary computational cost. In parallel, there are several AR-based attempts. They either rely on large DMs for final generation [35, 18], are limited to low resolution with additional up-sampling modules [9], or stitch perspective images without correct panorama geometry [77]. Overall, existing methods do not provide a unified and theoretically sound solution for conditional panorama generation, which we rethink through MAR modeling.
Autoregressive Modeling. Motivated by the success of large language models (LLMs) [4, 56], AR models, with great scalability and generalizability, are becoming a strong challenger to diffusion models. Several studies delve into visual AR with two distinct pipelines. One pipeline adopts raster-scan modeling [51, 45, 14, 15, 69, 60, 64, 27, 38], which treats the image as visual sentences and sequentially predicts the next token. However, this strategy does not fully exploit the spatial relationships of images. To make up for the shortcomings of raster-scan modeling, some works [6, 23, 70, 57, 29, 13] extend the concept of masked generative models [7] to AR modeling, namely MAR. Instead of predicting one token within one forward, the model randomly selects several candidate positions and predicts attending to the decoded positions. Intuitively, this modeling provides flexible control over arbitrary spatial positions simply by adjusting the order of tokens to be decoded. Hence, we resort to MAR as a flexible way to unify text and image-conditioned panorama generation while avoiding the theoretical defects of panorama DMs.
3 Method
This section is organized into three parts. First, we introduce the preliminaries of panorama projection and review the standard conditional AR modeling. Second, we analyze the disadvantages of this approach for unified conditional generation. Last, we describe our PAR in the following sections, including vanilla MAR architecture, consistency alignment, and circular padding.
3.1 Preliminaries
Equirectangular Projection is a common parameterization for panoramic images. It establishes a bijection between spherical coordinates $(\theta,\phi)$ and 2D image coordinates $(u,v)$, where $\theta\in[0,\pi]$ denotes the polar angle (latitude) and $\phi\in(-\pi,\pi]$ represents the azimuthal angle (longitude). This projection preserves complete spherical visual information while maintaining compatibility with conventional 2D convolutional operations. Given a 3D point $\mathbf{P}=(x,y,z)$ on the unit sphere $\mathbb{S}^{2}$, we first compute its spherical coordinates and then obtain ERP coordinates $(u,v)$ via linear mapping:
$\phi=\arctan\left(\frac{y}{x}\right),\quad\theta=\arccos(z),\quad u=\frac{\phi+\pi}{2\pi}W,\quad v=\frac{\theta}{\pi}H,$ (1)
where $W\times H$ denotes the target image resolution.
Conditional Autoregressive Models formulate conditional data generation as a sequential process through next token prediction. Given an ordered sequence of tokens $\{{x}_{t}\}_{t=1}^{T}$, the joint distribution $p({x|c})$ factorizes into a product of conditional probabilities:
$p({x|c})=\prod_{t=1}^{T}p({x}_{t}|{x}_{<t},c),$ (2)
where ${x}_{<t}\equiv({x}_{1},...,{x}_{t-1})$ denotes the historical context and $c$ indicates the conditions. This chain rule decomposition enables tractable maximum likelihood estimation, with the model parameters $\theta$ optimized to maximize:
$\mathcal{L}(\theta)=\mathbb{E}_{{x}\sim\mathcal{D}}\left[\sum_{t=1}^{T}\log p_{\theta}({x}_{t}|{x}_{<t},c)\right],$ (3)
where $\mathcal{D}$ represents the data distribution.
3.2 Rethinking Conditional Panoramic Image Generation
While DMs have demonstrated remarkable success in various image generation tasks, they rely on i.i.d. Gaussian noise assumption, which is violated by the spatial distortions introduced by the ERP transformation in Eq. 1, making diffusion-based methods inherently unsuitable for modeling panoramas. In this work, we therefore rethink conditional panorama generation from AR modeling, which is not constrained by such assumptions.
Building upon the standard conditional AR formulation in Eq. 2, we analyze text-to-panorama and panorama outpainting scenarios. The former takes text embeddings as $c$ while the latter takes observed image patches and optional textual prompts as conditions. In particular, this reformulates the two problems as follows:
$p({x}|{c})=\prod_{t\in\mathcal{S}_{u}}p({x}_{t}|{x}_{\mathcal{S}(t)},{c}),\quad\mathcal{S}(t)\subseteq(\mathcal{S}_{k}\cup\{1,...,t-1\}),$ (4)
where $\mathcal{S}_{k}$ denotes known token positions from the condition and $\mathcal{S}_{u}=\{1,...,T\}\setminus\mathcal{S}_{k}$ are targets. The context set $\mathcal{S}(t)$ dynamically aggregates:
$\mathcal{S}(t)=\underbrace{\{j|j\in\mathcal{S}_{k}\}}_{\text{condition}}\cup\underbrace{\{j|j<t,j\in\mathcal{S}_{u}\}}_{\text{generated content}},$ (5)
This formulation generalizes text-to-panorama ($\mathcal{S}_{k}=\emptyset$) and outpainting ($\mathcal{S}_{k}\neq\emptyset$) scenarios. However, traditional raster-scan modeling is not competent for the panorama outpainting task, as $\mathcal{S}_{k}$ is not necessarily at the beginning of the sequence.
Figure 2: Method illustration. (a) PAR utilizes a transformer $f$ to predict regions obscured by mask $M$, then employs these predictions as conditions to drive an MLP $\epsilon_{\theta}$ in generating continuous tokens. The VAE encoder $\mathcal{E}$ and decoder $\mathcal{D}$ are frozen. The dashed line indicates the inference phase. $c$ and $t$ denote textual embeddings and time-steps, respectively. (b) Both original and $\mathcal{T}_{v}$-augmented triples $(x,\epsilon,M)$ are processed by the same model, and then aligned through a consistency loss.
3.3 Masked Autoregressive Modeling with Consistency Alignment
We aim to learn a generalized AR model that naturally unifies text and image conditions and adapts to panoramic characteristics. An intuitive solution is MAR modeling, which allows generation in arbitrary order by marking the position of each token. However, existing foundation models are typically pre-trained on perspective images and cannot cope with panoramic scenarios. We delve into this problem and use the cyclic translation equivariance of ERP panoramas.
Vanilla Architecture. We start with a vanilla text-conditioned MAR model [13] to learn basic image generation without special constraints. Specifically, we use continuous tokens instead of discrete ones to reduce quantization error.
As illustrated in Fig. 2 (a), panoramic images $I$ are first compressed into latent representations $x$ using a VAE encoder $\mathcal{E}$. These representations are fed into a transformer model $f$, which adopts an encoder-decoder architecture. Within the encoder, $x$ is divided into a sequence of visual tokens via patchification. A subset of these tokens is randomly masked by replacing the corresponding regions with a designated [MASK] token. Sine-cosine positional encodings (PEs) are added to the token sequence to preserve spatial information.
Simultaneously, textual prompts are processed by a pre-trained text encoder [25], which encodes them into textual embeddings $c\in\mathbb{R}^{d}$, where $d$ matches the dimensionality of the visual tokens. The unmasked visual tokens are fused with the textual embeddings $c$ in the encoder, and the resulting contextualized representations are then passed to the decoder, where they are integrated with the masked tokens to facilitate predictive reconstruction.
Unlike the standard AR modeling, the decoder’s output is not used as the final result, but as a conditional signal $z$ to drive a lightweight denoising network, which is implemented by a three-layer multilayer perceptron (MLP) $\epsilon_{\theta}$. During the training phase, the noise-corrupted $x_{t}$ is denoised by $\epsilon_{\theta}$ under $z$ as follows:
$\mathcal{L}_{va}=\mathbb{E}_{t,\epsilon}\left[M\circ||\epsilon_{t}-\epsilon_{\theta}(x_{t}|t,z)||^{2}\right].$ (6)
In Eq. 6, $\circ$ is pixel-wise multiplication with binary mask $M$, where the value corresponding to the masked token is 1. The encoder-decoder $f$ and $\epsilon_{\theta}$ are jointly optimized.
Cyclic Translation Consistency. Current foundation models [46, 13] are predominantly trained on perspective images, leading to suboptimal feature representations for ERP-formatted panoramas. We introduce a cyclic translation operator $\mathcal{T}_{v}$ to address this domain gap and construct semantically equivalent panorama pairs, where $v$ indicates arbitrary translation distance. As shown in Fig. 2 (b), we construct an input pair, $(x,\epsilon,M)$ and $(x^{\prime},\epsilon^{\prime},M^{\prime})=(\mathcal{T}_{v}(x),\mathcal{T}_{v}(\epsilon),\mathcal{T}_{v}(M))$. Both inputs share identical PE maps to prevent trivial solutions. The main model processes these pairs to produce outputs $y$ and $y^{\prime}$, respectively. We enforce equivariance by aligning $\mathcal{T}_{v}(y)$ with $y^{\prime}$ only on masked regions through the objective:
$\mathcal{L}_{consistency}=M^{\prime}\circ||\mathcal{T}_{v}(\epsilon_{\theta}(x_{t}|t,x,M))-\epsilon_{\theta}(x_{t}^{\prime}|t,x^{\prime},M^{\prime})||^{2},$ (7)
The final learning objective can be written as follows:
$\mathcal{L}=\mathcal{L}_{va}+\lambda\mathcal{L}_{consistency},$ (8)
where $\lambda$ is set as 0.1 in our experiments.
Discussion. The equivariance constraint holds exclusively for ERP panoramas due to their inherent horizontal continuity. For perspective images, applying $\mathcal{T}_{v}$ artificially creates discontinuous boundaries in the center of the shifted image, which violates two critical premises. 1) Semantic equivalence. Unlike panoramas, shifted perspective images lose semantic consistency due to non-periodic scene structures. 2) Model prior. Generation models inherently enforce feature continuity in image interiors, making it infeasible to produce fragmented content.
Circular Padding. While the AR model incorporates panoramic priors, the VAE’s latent space processing may still introduce discontinuities. Standard convolution operations in VAE provide sufficient context for central regions but suffer from information asymmetry at image edges - pixels near boundaries receive unidirectional receptive fields. To address this issue, we propose a dual-space circular padding scheme: pre-padding and post-padding. The former is applied in pixel space for the VAE encoder, while the latter is used in latent space for the VAE decoder. Specifically, we crop a particular portion of the left and right regions and concatenate them to the other side, as formulated in Eq. 9.
$C_{r}(x)=\mathrm{concat}(x[...,-r*W/2],x,x[...,r*W/2]),$ (9)
where $x\in\mathbb{R}^{H\times W}$ and $C_{r}(x)\in\mathbb{R}^{H\times(1+r)W}$, indicate the original and the padded tensor, respectively. $r$ indicates the ratio of padding width to that of the original images or latent variables. After the VAE transformation, the edge positions have sufficient context information. Therefore, the areas that belong to the padding are discarded.
4 Experiment
Figure 3: Visual comparisons with previous methods on text-to-panorama task. Previous methods neglect the circular consistency characteristics (yellow box), or suffer from repetitive objects (red box), artifact generation (blue box), and inconsistent combinations (green box). The highlighted portions in captions indicate text-image alignment failure cases of baseline methods.
Table 1: Performance on text-to-panorama task. PO and PE represent panorama outpainting and editing, respectively. Bold and underline indicate the first and second best entries.
Modeling Method ##\##params T2P PO PE FAED $\downarrow$ FID $\downarrow$ CS $\uparrow$ DS $\downarrow$
Diffusion PanFusion [71] 1.9B - - 5.12 45.21 31.07 0.71
AR Text2Light [9] 0.8B - - 68.90 70.42 27.90 7.55
PanoLlama [77] 0.8B - - 33.15 103.51 32.54 13.99
PAR (ours) 0.3B 3.39 41.15 30.21 0.58
PAR (ours) 0.6B 3.34 39.31 30.34 0.57
PAR (ours) 1.4B 3.75 37.37 30.41 0.58
4.1 Experiment Settings
Implementation Details. We utilize NOVA [13] as the initialization and set the resolution as $512\times 1024$. Our model is trained for 20K iterations with a batch size 32. We employ an AdamW optimizer [34]. The learning rate is $5\times 10^{-5}$ with the linear scheduler. In the inference stage, we set the CFG [22] coefficient as 5. In this paper, the masking sequence for training and the sampling sequence for inference are initialized with a uniform distribution unless otherwise stated.
Datasets and Metrics. We mainly use Matterport3D [5] for comparisons. The split of the training and validation set follows PanFusion [71]. We use Janus-Pro-7B [8] to generate the captions.
Several metrics are used in this paper. Fréchet Inception Distance (FID) [20] measures the similarity between the distribution of real and generated images in a feature space derived from a pre-trained inception network. Lower FID scores indicate better realism of images. CLIP Score (CS) [44] evaluates the alignment between text and image. Furthermore, as FID relies on an Inception network [52] trained on perspective images, we also report the Fréchet Auto-Encoder Distance (FAED) [39] score, which is a variant of FID customized for panorama. To measure the cycle consistency of panoramic images, we adopt Discontinuity Score (DS) [10].
Baselines. Since only a few approaches combine T2P and PO, we compare PAR with specialist methods separately. Specifically, for T2P, we compare with several diffusion-based and AR-based methods, including PanFusion [71], Text2Light [9], and PanoLlama [77]. We take AOG-Net [35] and 2S-ODIS [37] as PO baselines. All experiments are implemented on the Matterport3D dataset.
Figure 4: Qualitative comparisons of panorama outpainting on the Matterport3D dataset. PAR-1.4B is used for this task, where PAR w/o prompt means the textual prompt is set as empty.
Table 2: Panorama outpainting results on the Matterport3D dataset. For FID-h, we horizontally sample 8 perspective images with FoV=$90\degree$ for each panorama and then calculate their FID.
Method T2P PO PE FID $\downarrow$ FID-h $\downarrow$
AOG-Net [35] - - 83.02 37.88
2S-ODIS [37] - - 52.59 35.18
PAR w/o prompt 41.63 25.97
PAR w/ prompt 32.68 12.20
4.2 Main Results
Text-to-Panorama. Tab. 1 shows the quantitative comparison results. Within the scope of AR-based methods, our PAR significantly outperforms previous approaches regarding generation quality and continuity score. Our method is slightly inferior in the CS metric, probably because CLIP [44] is pre-trained on massive perspective images, and panoramic geometry is not well aligned with perspective images. Moreover, our model performs decently compared with a strong diffusion-based baseline, PanFusion, with only 0.3B parameters. Specifically, PAR-0.3B outperforms PanFusion with 1.73 and 4.06 on FAED and FID, with a discontinuity score of 0.58. Scaling up the parameters helps achieve better results, which is further illustrated in Sec. 4.3.
Fig. 3 provides visual comparisons with the aforementioned baseline methods, including diffusion-based and AR-based models. Our model shows better instruction-following ability, more consistent generation results, and avoids duplicate generation. Specifically, Text2Light cannot obtain satisfactory synthesized results. PanoLlama cannot understand panorama geometry and neglects the characteristic of cycle consistency. PanFusion produces more reasonable results, but sometimes faces artifact problems. For example, the first and second columns of the third row show a room without an exit.
Panorama Outpainting. Tab. 2 compares our method with AOG-Net and 2S-ODIS. We use the model trained with Eq. 8 and inference following Eq. 4, where $\mathcal{S}_{k}$ indicates the positions within the mask. Our model achieves 32.68 and 12.20 on FID and FID-h, respectively. FID-h is similar to FID, but the difference is that it extracts eight perspective views in the horizontal direction with a 90-degree FoV for each panoramic image, which pays more attention to the generation quality of local regions. Fig. 4 visualizes the synthesized results. Existing methods suffer from local distortion (the $3^{rd}$ column) or counterfactual repetitive structures (the $2^{nd}$ column). Our method overcomes these problems and still produces reasonable results without textual prompts.
4.3 Ablation Study
Figure 5: Scaling parameters and training compute improve fidelity and soundness. Two cases are shown with 3 model sizes and 3 different training stages. From top to bottom: 0.3B, 0.6B, 1.4B. From left to right: $25\%$, $50\%$, $100\%$ of the training process.
Figure 6: Panorama outpainting on OOD dataset. The images are from SUN360, which is out of the distribution of our training data. PAR generates realistic panorama images while previous methods have problems like artifacts, or unrealistic results.
Figure 7: Panoramic image editing. We design several challenging textual prompts to enforce the model to regenerate the contents outside the red box while maintaining the inside contents. Note that our PAR performs zero-shot task generalization.
Figure 8: Left: The curve of DS about $r_{post}$. Right: Visualizations of different circular padding strategies. Each part consists of three images, where the top one represents the panoramic image generation result, the left and right ones represent the edge stitching results, respectively. The red box indicates the semantic discontinuity, and the green box highlights the seam after post-padding.
Scalability. We train with three different parameter sizes, namely 0.3B, 0.6B, and 1.4B, respectively. In Tab. 1, FID and CS improve with increasing parameters. We also visualize the performance under different training stages, including $25\%$, $50\%$, and $100\%$ of the whole training process. As shown in Fig. 5, the improvement of fidelity and soundness is consistent with the scaling of parameters and computation. Larger models learn data distributions better, such as details and panoramic geometry.
Generalization. Fig. 6 visualizes the outpainting results of PAR on OOD data. SUN360 [63] is used for evaluation, which has different data distributions from Matterport3D. PAR shows decent performance across various scenarios. We also compared with several PO baselines trained on the SUN360 dataset. However, Diffusion360 [17] suffers from realistic scenes and lacks details. Panodiff [59] also encounters artifact problems (red sky in the $1^{st}$ row).
PAR can also adapt to different tasks, such as panoramic image editing. As shown in Fig. 7, the model needs to preserve the content inside the mask and regenerate according to new textual prompts. We designed three challenging text prompts, whose content differs from the training datasets, and contains specific conflicts with the original image. For example, fire vs. yard in (b), snow vs. current season of scene in (c), and ocean vs. current place of the scene in (d). Interestingly, our model can generate them well and understands panoramic characteristics. As shown in Fig. 7 (b), the left and right sides of the image can be stitched to form a whole fire.
Table 3: Ablations on cyclic consistency loss.
Method FID $\downarrow$ CS $\uparrow$ DS $\downarrow$
w/o $\mathcal{L}_{consistency}$ 39.55 30.25 0.57
w/ $\mathcal{L}_{consistency}$ 37.37 30.41 0.58
Translation Equivariance. Tab. 3 demonstrates the effectiveness of consistency loss. Eq. 7 helps the model adapt to panoramic characteristics, thus improving generation quality. Fig. 8 (left) shows the convergence of $r$. Post-padding significantly improves DS when $r_{post}>0$ as it helps to smooth stitching seams in the pixel space. However, the model without pre-padding suffers from higher discontinuity, indicating the misalignment between the training and testing phases. With the same $r_{pre}$, DS converges quickly when $r_{post}>0.125$. A similar phenomenon is observed with $r_{post}$ fixed. Specifically, the curve of $r_{pre}=0.25$ and $r_{pre}=0.5$ almost coincide.
Fig. 8 (right) visually explains the difference between the two padding strategies. When $r_{pre}=0$, semantic inconsistency exists between the left and right edges of the panoramic image. Post-padding cannot repair it even with a large padding ratio, and the green box bounds the smoothed seam. In contrast, the model generates consistent contents when pre-padding is applied.
5 Conclusion
This paper proposes PAR, a MAR-based architecture unifying text and image-conditioned panoramic image generation. We delve into the cyclic translation equivariance of ERP panoramas and correspondingly propose a consistency alignment and circular padding strategy. Experiments demonstrate PAR’s effectiveness on text-to-panorama and panorama outpainting tasks, whilst showing promising scalability and generalization capability. We hope our research provides a new solution for the panorama generation community.